# EVT: Efficient View Transformation for Multi-Modal 3D Object Detection

### v2

@@ 3.1 Adaptive Sampling and Adaptive Projection
Table 1: Performance comparison on the nuScenes validation and test sets. The results are obtained without model ensemble or test-time augmentation. ‘L’ and ‘C’ denote LiDAR and Camera, respectively.
3.2 Group-wise Query Initialization
The proposed group-wise query initialization method effectively generates queries for transformer-based detection frameworks. First, group-wise heatmaps are predicted from the multi-modal BEV feature map $\mathbf{BEV}_{\text{fuse}}$. The groups are predefined, with each group containing similarly-sized classes. The predicted heatmaps have scores ranging from 0 to 1, representing the likelihood that each BEV pixel corresponds to the center of an object. The heatmap head is supervised by 2D Gaussian distributions centered at each object’s location. Next, the top-k keypoints are selected from each heatmap group. The positions of these keypoints are used as reference points for queries in BEV space.
This method initializes each query feature using group-wise learnable parameters. Queries within the same group share these parameters, which enables them to effectively capture the characteristics of group. This demonstrates better performance in the multi-layer transformer decoder than traditional query initialization approaches that sample features directly from the input feature map [47, 1]. Further details are provided in Sec. 4.4.
3.3 Enhanced Query Update Framework
For more effective query updates within a transformer decoder based on deformable attention [49], we propose the corner-aware sampling and position-embedded feature mixing methods.
Corner-aware Sampling.
The proposed corner-aware sampling method shifts the sampling reference from the center of the object to each corner and rotates sampling offsets according to the object’s heading. This approach addresses the limitations of existing center-based sampling methods, which generate sampling offsets around the object’s center, resulting in difficulties capturing the full spatial extent and shape of the objects.
First, the initial points $\{(\Delta x_{i}^{\prime},\Delta y_{i}^{\prime})\}$ are generated from the query feature $\mathbf{q}$ using a linear layer:
where $N_{\text{p}}$ denotes the number of sampling points.
Subsequently, the sampling offsets $\{(\Delta x_{i},\Delta y_{i})\}$ and sampling points $\{(x_{i},y_{i})\}$ are determined via a geometric transformation as follows:
where $(I_{j},I^{\prime}_{j})\,\in\{(1,1),(1,-1),(-1,1),(-1,-1)\}$. $j$ denotes the index of each corner and is the remainder of $i$ divided by four, the number of corners. $(x_{c},y_{c})$ denote the coordinates of the object’s center. $l$, $w$, and $\theta$ denote the length, width, and yaw of the predicted bounding box, respectively, obtained from the regression head of the previous transformer decoder layer. In the first decoder layer, where there is no preceding layer to provide bounding boxes, the equation is satisfied by setting $l$, $w$, and $\theta$ to zero.
The features are sampled by bilinear interpolation at each sampling point on the multi-modal fused features $\mathbf{BEV}_{\text{fuse}}$ as follows:
$g_{i}={\rm B}(\mathbf{BEV}_{\text{fuse}},(x_{i},y_{i}))\in\mathbb{R}^{C},$ (12)
where ${\rm B}(\cdot)$ denotes the bilinear interpolation. $g_{i}$ represents the sampled feature that captures important key tokens by leveraging the geometric properties of the query’s bounding box.
Position-embedded Feature Mixing.
Accurate locations of the sampled features are essential for providing a more detailed representation of the bounding boxes. In AdaMixer [9], the adaptive channel and spatial mixing are combined to leverage query guidance and effectively decode sampled features. However, this method is less effective for decoding the outputs of the corner-aware sampling, as shown in Tab. 3 (e), which involves complex geometric transformations of the sampling points. To address this, we propose position-embedded feature mixing, an enhancement based on AdaMixer that improves the decoding capability by incorporating embeddings for sampling offsets.
First, the sampling offsets are embedded as position vectors using sinusoidal position encoding [33], followed by a linear layer. Then, the position-embedded sampled features $G\in\mathbb{R}^{N_{\text{p}}\times C}$ are obtained by element-wise addition of the sampled feature $g_{i}$ and the position vector $e_{i}$:
$e_{i}$ $={\rm Linear}(\Phi_{pos}((x_{i},y_{i})))\in\mathbb{R}^{C}$ (13)
$G$ =$=\{(g_{i}+e_{i})\mid i\in{0,1,...,N_{\text{p}}-1}\},$ (14)
where $\Phi_{pos}(\cdot)$ denotes sinusoidal position encoding.
$W_{c}$ $={\rm Linear}(\mathbf{q})\in\mathbb{R}^{C\times C}$ (15)
$G_{c}$ $={\rm ReLU}({\rm LN}(G\times W_{c})).$ (16)
$G_{cs}$ $={\rm ReLU}({\rm LN}(G_{c}^{T}\times W_{s})).$ (18)
Finally, the query feature is updated as follows:
$\mathbf{q^{\prime}}=\mathbf{q}+{\rm Linear}({\rm Flatten}(G_{cs})),$ (19)
where $\mathbf{q^{\prime}}$ represents the updated query feature that is enhanced for a more accurate feature representation by incorporating the positional information of each sampled point.
4 Experiments
@@
4.1 Implementation Details
For the image backbone, either ResNet-50 [10] with a resolution of $704\times 256$ or V2-99 [15] with $1600\times 640$ is used with the feature pyramid network (FPN) [22]. For the LiDAR backbone, VoxelNet [46] is used with an ROI of $[-54.0m,54.0m]$ along $(X,Y)$ and $[-5.0m,3.0m]$ along $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$. In ASAP, we use 4 points for each grid cell and the last two stages of each image backbone for view transformation. The size of the unified BEV feature map size is $180\times 180$. The object categories are divided into 6 groups, with 150 queries per group, resulting in a total of 900 queries. The transformer decoder has 6 layers, and the corner-aware sampling generates 16 sampling points per query. The feature dimension is set to 256.
@@ 4.3 Comparison with State-of-the-art Methods
As shown in Tab. 1, we compare EVT and its LiDAR-only model, EVT-L, with existing methods on the nuScenes validation and test sets. The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, outperforming all existing state-of-the-art approaches on both the nuScenes validation and test sets. In particular, it surpasses recent methods, such as UniTR [35] by 0.8% NDS and 1.7% mAP, MSMDFusion [14] by 1.3% NDS and 1.1% mAP, and SparseFusion [39] by 1.5% NDS and 0.6% mAP. Furthermore, EVT shows a performance improvement of 3.2% NDS and 4.9% mAP compared with the LiDAR-only model EVT-L. In contrast, TransFusion [1] shows only a 1.5% NDS and 3.4% mAP improvement over its LiDAR-only model, TransFusion-L. This indicates that EVT effectively utilizes camera data through the proposed view transformation method.
@@
Figure 6: Comparison of query initialization strategies on the nuScenes validation set. For the multi-layer transformer decoder, group-wise query initialization without feature sampling achieves the best performance.
Attention Reference scale rotate feature mixing position-embedded NDS mAP
@@
Table 3: Comparison of the cross-attention methods within the transformer decoder. The proposed corner-aware sampling and position-embedded feature mixing achieve significant performance improvements.
@@
4.4 Ablation Studies
In this section, we describe the validation of each component of the proposed method on the nuScenes validation set. Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [48] and the query denoising strategy [16].
Adaptive Sampling and Adaptive Projection
Tab. 2 displays the results of the ablation conducted on ASAP by retraining the entire model. In these experiments, ResNet [10] is used as the backbone network with a resolution of $704\times 256$. (a) shows the performance of EVT-L, the LiDAR-only model. In (b), the vanilla view transformation (VT) method, without LiDAR guidance, employs a 3D-to-2D projection of predefined 3D points for feature sampling.
In (c), AS achieves improvements of 0.8% NDS and 1.5% mAP compared to the vanilla VT method. The sampling points of the vanilla VT and AS are visualized in Fig. 4. The sampling points of AS are adaptively generated in the ROIs.
In (d), the entire ASAP achieves improvements of 1.0% NDS and 1.6% mAP compared to the vanilla VT method (b). The BEV feature maps of each component are visualized in Fig. 5. ASAP effectively transforms image features into BEV space and resolves ray-directional misalignment.
Group-wise Query Initialization
Different query initialization methods using the heatmap head are ablated in Fig. 6. All methods utilize a top-K keypoints selection approach to define queries, but differ in their specific implementations. (a) directly samples features from the input BEV feature map and combines them with group-wise embedding, which is similar to the approach commonly used in 3D object detectors [47, 1]. (b) initializes queries with group-wise embedding using global top-K selection. (c) is our proposed group-wise query initialization strategy, which initializes queries using top-K selection and embedding for each group.
In the results, the direct feature sampling approach in (a) is more effective in a single-layer transformer decoder, outperforming (c) by 0.6% NDS. However, in a multi-layer transformer decoder, this leads to performance degradation. Specifically, with 6 layers, (c) surpasses (a) by 1.0% NDS. Additionally, group-wise top-K (c) consistently outperforms global top-K (b) across all layers by approximately 0.1% NDS.
Enhanced Query Update Framework
We ablate the corner-aware sampling method, as demonstrated in Tab. 3 (a)-(d). (a) is the standard multi-head cross attention [33], and (b)-(c) are the standard deformable attention [49]. Specifically, in (c), the sampling offsets are scaled according to the bounding box size and rotated according to the heading. (d) is the proposed corner-aware sampling method, which samples features from the bounding box corners of each query and rotates them according to the heading. The corner-aware sampling method (d) achieves improvements of 0.8% NDS and 0.5% mAP compared to method (c).
The results of the ablation study on the feature mixing method are displayed in Tab. 3 (e)-(f). In (e), the standard feature mixing method of AdaMixer [9] does not yield any improvements over our method. In (f), the proposed position-embedded feature mixing demonstrates improvements of 0.4% NDS and 0.7% mAP compared to (d).
@@ 4.4 Ablation Studies
As described in Sec. 3.3, the corner-aware sampling is applied starting from the second transformer decoder layer. We experimented on two different approaches to incorporate the corner-aware sampling into all transformer decoder layers. The first approach involves adding regression heads during the query initialization stage (Sec. 3.2) to predict the bounding boxes, whereas the second approach uses learnable bounding boxes from SparseBEV [23]. However, as shown in Tab. 4, neither of these approaches led to any noticeable performance improvement. Furthermore, the experiment using learnable bounding boxes results in performance degradation.
@@ 5 Conclusion
In this paper, we have proposed EVT, a novel transformer-based 3D object detector. The proposed view transformation method, ASAP, enables robust and efficient LiDAR-camera fusion by leveraging LiDAR guidance. In addition, the proposed group-wise query initialization method provides effective queries for a multi-layer transformer decoder. The corner-aware sampling and position-embedded feature mixing methods effectively utilize geometric properties to update queries, and these methods can easily be extended to other models. We expect EVT to provide valuable insights into multi-modal 3D object detection.

### v3

@@ 4 Experiments
Table 1: Performance comparison on the nuScenes validation and test sets. The results are obtained without model ensemble or test-time augmentation. ‘L’, ‘C’ and ‘T’ denote LiDAR, camera and temporal fusion, respectively.
4.1 Implementation Details
For the image backbone, we use ResNet-50 [11] with a resolution of $704\times 256$ or V2-99 [18] with $1600\times 640$ with FPN [26]. The LiDAR backbone is VoxelNet [57] with an ROI of $[-54.0m,54.0m]$ in $(X,Y)$ and $[-5.0m,3.0m]$ in $Z$, with a voxel size of $(0.075m,0.075m,0.2m)$. In ASAP, we use four sampling points for each grid cell. The multi-modal BEV feature map size is $180\times 180$. Following [52], object groups in 3.2 are defined as: (1) car, (2) truck, construction vehicle, (3) bus, trailer, (4) barrier, (5) motorcycle, bicycle, (6) pedestrian, traffic cone. Each group contains 150 queries, resulting in a total of 900 queries, and the corner-aware sampling generates 16 sampling points per query, both of which were empirically determined. The transformer decoder has six layers, and the feature dimension is set to 256.
@@ 4.3 Comparison with State-of-the-art Methods
As shown in Tab. 1, we compare EVT and its LiDAR-only model, EVT-L, with existing methods on the nuScenes validation and test sets. The multi-modal EVT achieves 75.3% NDS and 72.6% mAP, surpassing all previous approaches on both the nuScenes validation and test sets. In particular, it surpasses recent methods, such as UniTR [40] by 0.8% NDS and 1.7% mAP, MSMDFusion [17] by 1.3% NDS and 1.1% mAP, and SparseFusion [45] by 1.5% NDS and 0.6% mAP. Furthermore, EVT shows a performance improvement of 3.2% NDS and 4.9% mAP compared with the LiDAR-only model EVT-L. In contrast, TransFusion [1] shows only a 1.5% NDS and 3.4% mAP improvement over its LiDAR-only model, TransFusion-L. This indicates that EVT effectively utilizes camera data through the proposed view transformation method.
@@
Method # Decoder Layers & NDS (%)
1 2 3 4 5 6
(a) Learnable Init. 56.8 64.4 67.5 68.7 68.9 69.6
(b) Heatmap Init. 70.4 70.5 70.5 70.3 70.5 70.7
(c) Mixed init. 69.2 70.3 70.5 70.5 70.7 71.1
+ Group-wise 69.8 71.1 71.3 71.2 71.4 71.7
Table 3: Comparison of query initialization strategies on the nuScenes validation set. (a) Fully learnable initialization, (b) Fully heatmap-based initialization, (c) Our mixed initialization strategy: the first row denotes instance-wise mixed query selection and the second row denotes group-wise mixed query selection.
Attention Reference scale rotate feature mixing position-aware NDS mAP
@@
Table 4: Comparison of attention methods within the transformer decoder. The proposed geometry-aware cross-attention, which includes corner-aware sampling and position-aware feature mixing, achieves significant performance improvements.
Method NDS mAP
ResNet-50 baseline (24 epochs) 47.8 37.2
+ geometry-aware cross-attention 49.1 37.8
ResNet-50 baseline (90 epochs) 53.5 42.7
+ geometry-aware cross-attention 54.7 43.4
Table 5: Impact of geometry-aware cross-attention on StreamPETR [42]. All models are trained in our experiments.
4.4 Ablation Studies
In this section, we describe the validation of each component of the proposed method on the nuScenes validation set. Unless otherwise specified, all experiments are conducted using the proposed LiDAR-only model EVT-L, trained for 10 epochs with CBGS [59] and the denoising strategy [19].
Adaptive Sampling and Adaptive Projection
In Tab. 2, the ablation results for ASAP are obtained by retraining the entire model. In these experiments, ResNet [11] is used as the backbone network with a resolution of $704\times 256$. (a) shows the performance of EVT-L, the LiDAR-only model. In (b), the vanilla view transformation (VT) method, without LiDAR guidance, employs a 3D-to-2D projection of predefined 3D points for feature sampling.
In (c), AS achieves improvements of 0.8% NDS and 1.5% mAP compared to the vanilla VT method. The sampling points of the vanilla VT and AS are visualized in Fig. 4. The sampling points of AS are adaptively generated in highly object-relevant regions of the image.
In (d), the entire ASAP achieves improvements of 1.4% NDS and 2.0% mAP compared to (b), while maintaining high efficiency with only a 3ms latency increase. The BEV feature maps of each component are visualized in Fig. 5. ASAP effectively transforms image features into BEV space and resolves ray-directional misalignment.
Group-wise Mixed Query Selection
As shown in Tab. 3, we conduct an ablation study on query initialization strategies. (a) fully learnable initialization, where both features and positions are learnable [29, 30, 27, 46, 12]. (b) fully heatmap-based initialization, where positions are obtained from high-score heatmap keypoints, and features are sampled at those locations [58, 1, 51, 55, 44]. (c) our mixed initialization strategy, which combines heatmap-derived positions with either instance-wise or group-wise embeddings.
These results highlight the importance of query initialization. Notably, fully learnable initialization without any prior information consistently yields the lowest performance. While (b) achieves the best performance in a single-layer transformer decoder, (c) outperforms all other strategies in multi-layer settings, improving NDS by over 1% at the last layer. Additionally, group-wise embeddings allow each group to learn more generalized feature representations, leading to meaningful improvements.
Geometry-aware Cross-Attention
We ablate the corner-aware sampling method, as shown in Tab. 4 (a)-(d). (a) employs multi-head attention [38], and (b)-(c) use deformable attention [60]. Specifically, in (c), the sampling offsets are scaled based on the bounding box size and rotated according to the heading. (d), the proposed corner-aware sampling method, samples features from bounding box corners aligned with the heading, achieving improvements of 0.8% NDS and 0.5% mAP compared to (c).
In (e) and (f), AdaMixer [10] fails to improve performance, as it does not effectively preserve the structured sampling introduced by corner-aware sampling. In contrast, our position-aware feature mixing explicitly incorporates positional embeddings, leading to improvements of 0.4% NDS and 0.7% mAP compared to (d). As a result, our proposed geometry-aware cross-attention achieves overall improvements of 1.2% NDS and 1.2% mAP.
Additionally, as shown in Tab. 5, we validate the applicability of geometry-aware cross-attention by integrating it into a camera-only 3D detector. Modifying only the cross-attention layers in StreamPETR [42] improves performance without further adjustments.
@@ 4.4 Ablation Studies
As described in Sec. 3.3, the corner-aware sampling method is applied starting from the second transformer layer. We explore two different approaches to extend the corner-aware sampling to all transformer layers. The first approach involves adding regression heads during query initialization in Sec. 3.2 to predict bounding boxes, whereas the second approach uses learnable bounding boxes. However, as shown in Tab. 6, neither approach resulted in any noticeable performance gains, and the experiment with learnable bounding boxes even led to performance degradation.
@@ 5 Conclusion
We propose EVT, a novel multi-modal 3D object detector based on BEV representation, enhancing both efficiency and accuracy. Our method introduces ASAP, an efficient LiDAR-camera fusion method that leverages LiDAR guidance for accurate view transformation. Additionally, the proposed group-wise mixed query selection improves initial feature representation through shared embeddings. The geometry-aware cross-attention refines queries using geometric properties and can be easily extended to other models. We expect EVT to provide valuable insights into multi-modal 3D object detection.